Tests on the breaking strengths if 10 cables were made and showed a mean strength of 1200 psf and a standard deviation of 100 psf. Estimate the (a) mean and the (b) standard deviation of the population of all the cables being produced by the company.
1
Expert's answer
2020-05-14T17:41:02-0400

a)mean=μ=1200 pcf

b) "σ=\sqrt{s^2*(n-1)/n}= 
\sqrt{100^2*9/10}=94.868 pcf"